# Predictions for new physics in the CMB damping tail

### v1

@@ III.2 Early dark energy
We have computed the predictions from an scalar field EDE model following Refs. Poulin et al. (2019); Smith et al. (2020) using a scalar field potential which scales as $\phi^{6}$ around its minimum, thereby fixing the rate at which the EDE energy density redshifts after its peak to be $\rho_{\phi}\propto a^{-4.5}$ Turner (1983). In addition to this we have computed the predictions using a fluid parameterization Simon (2024); Lin et al. (2019); Poulin et al. (2023) in order to explore whether there are aspects of the small-scale CMB power spectrum in these models that are unique to the scalar field dynamics. In the fluid case we have left the equation of state after the peak EDE contribution as an additional free parameter.
The constraints for both these variations are shown in figure 2. The fluid approximation deviates slightly more than the axion-like early dark energy model. However, generally the two models behave very similarly. We do observe that they deviate especially in the temperature autocorrelation on the smallest scales (at $\ell\gtrsim 3000$), but in a way that is mostly compatible with the expected uncertainty of ACT DR6 and somewhat degenerate with the non-linear (baryonic feedback) treatment. Whether such a component would be measurable is not clear, though might be possible with CMB-S4 data and a tight control of the baryonic effects of structure formation from other observations as well as accurate and precise foreground removal.
@@
We show the constraints for this model in figure 3. For the original cold NEDE model the constraints are almost identical to that of the axion-like EDE, except for a clear upturn in the TE spectra around $\ell\gtrsim 1500$ that could be detectable with near-future data. On the other hand, tNEDE clearly shows a very different phenomenology on the smallest scales, with a non-zero posterior for the energy density of the trigger field today, $\Omega_{0,t}=0.0042^{+0.0018}_{-0.0030}$ and a mass $m_{t}=2.15509^{+0.4476}_{-0.5755}\times 10^{-27}\ {\rm eV}$. In this case the model is generally more consistent with $\Lambda$CDM due to the additional suppression of the small-scale power– $S_{8}=0.818^{+0.022}_{-0.019}$ in tNEDE versus $S_{8}=0.847\pm 0.013$ in NEDE. The two models show similar features in the TE spectrum around $\ell\sim 1000$ that might be detectable. We note that while the mean value in the temperature autocorrelation for this variation is much more discrepant from the $\Lambda$CDM reference due to the additional suppression of small-scale power, the $1\sigma$ contour is quite broad, making it somewhat uncertain how well this feature may be measured with future CMB data.
@@
It is interesting to note that, in the absence of an ultra-light scalar field, the predicted small-scale power spectra for all EDE models considered here are very similar, showing a clear upturn in both TT and EE starting around $\ell\gtrsim 3000$. As we argue in section IV, such a feature is generic for any model that addresses the Hubble tension by injecting a significant amount of energy density over a limited amount of time.
@@ III.3 Varying electron mass
There are several indications that our current standard model of particle physics is not the final description of the universe (lacking explanations for dark energy and dark matter as well, a neutrino mass mechanism, as well as unification with gravity at higher energies) and therefore a natural question is to ask what can happen if fundamental ‘constants’ are allowed to vary in time. In particular, an electron mass, $m_{e}$, which was slightly larger around recombination than it is today has proven to be surprisingly effective at resolving the cosmic conundrum of the Hubble tension as well as being slightly favored by recent DESI data, see Hart and Chluba (2020); Sekiguchi and Takahashi (2021); Schöneberg and Vacher (2025); Schöneberg et al. (2022). With a larger value of $m_{e}$ the redshift of recombination is increased, leading to a smaller sound horizon and therefore a larger allowed Hubble constant while keeping the CMB angular sound horizon scale unchanged. Indeed, in Schöneberg and Vacher (2025); Sekiguchi and Takahashi (2021) it has been argued that the CMB might be entirely insensitive to such a change at first order if other cosmological parameters are varied concurrently. See Hart and Chluba (2018); Chiang and Slosar (2018); Hart and Chluba (2022); Chluba and Hart (2023); Hart and Chluba (2020); Sekiguchi and Takahashi (2021); Lynch et al. (2024) for further constraints, and Schöneberg and Vacher (2025) for a review.
@@
Here we implement the model as described in Schöneberg and Vacher (2025) using a phenomenological approach – noting that a full UV-complete theory or even an effective field theory of such electron mass variations is still lacking. Despite this, see Ref. Schöneberg and Vacher (2025) for several proposals for a mechanism underlying such phenomenological variations. We describe the model using a single additional parameter $m_{e}^{\mathrm{early}}/m_{e}^{\mathrm{late}}$ describing the relative difference of the electron mass around CMB times compared to that measured in the laboratory today. We assume a transition in the dark ages where electromagnetic observational data is sparse, see also Ref. Schöneberg and Vacher (2025) for more details on other observational probes for this model, including a discussion on why BBN light element abundances are currently not constraining enough to restrict such variations.
@@
We show the predicted power spectra in figure 4. We observe that the varying $m_{e}$ model can leave all CMB scales invariant with respect to $\Lambda$CDM and is therefore mostly indistinguishable, even at small scales, despite having a much larger value of $H_{0}$ on average, around $H_{0}=71.2\pm 0.7$km/s/Mpc (with $m_{e}^{\mathrm{early}}/m_{e}^{\mathrm{late}}=1.0198\pm 0.0044$, a roughly 2% increase in the electron mass at CMB times) Schöneberg and Vacher (2025); Sekiguchi and Takahashi (2021). A similar result can be found when including curvature (green curve and colored region in figure 4). In this case there are larger oscillations in the mean prediction, but overall the result is not far outside the expected uncertainty bands for near-future CMB surveys and the mean value is still mostly compatible with zero. In this case we find $H_{0}=71.8\pm 1.1$km/s/Mpc with $m_{e}^{\mathrm{early}}/m_{e}^{\mathrm{late}}=1.037\pm 0.013$, a roughly 4% increase in the electron mass at CMB times, and $\Omega_{k}=(-6.5\pm 4.5)\cdot 10^{-3}$.
@@ IV Discussion
One part of the explanation is due to the lack of balance between the small-scale effect of $\theta_{\rm d}$ and $n_{s}$. As was shown in Ref. (Smith and Poulin, 2024, III.A), both parameters need to increase to keep the small-scale CMB invariant.
Their Eq. (7) reads
@@
This implies a clear degeneracy along which these parameters must move to keep the small-scale CMB relatively unchanged. Figure 7 shows the posteriors of $\theta_{\rm d}$ and $n_{s}$ for $\Lambda$CDM, WZDR, and axion-like EDE. The black line shows the degeneracy between these parameters which leaves the small-scale CMB relatively unchanged (compared to the $\Lambda$CDM mean values), as given in equation 1 with $\delta\theta_{\rm d}\equiv\theta_{\rm d}-\bar{\theta}^{\Lambda{\rm CDM}}_{\rm d}$ and $\delta n_{s}\equiv n_{s}-\bar{n}^{\Lambda{\rm CDM}}_{s}$. It is clear that the axion-like EDE model lies along this degeneracy whereas WZDR is significantly above it. Models which are above the line have a damping scale which is larger, leading to enhanced damping on small scales.
@@
Variations in the electron mass leave both the damping scale and $n_{s}$ unchanged. Additionally, as argued in Schöneberg and Vacher (2025) there is a strong degeneracy for all underlying scales (damping, angular, equality, …) involved in the CMB that should cause this model to be very close to $\Lambda$CDM, which is what we observe.
@@ V Conclusions
With the advent of precision cosmology the emerging Hubble tension has become a serious problem for the standard $\Lambda$CDM model. In particular since the arcminute-scale CMB anisotropies have been observed by the Planck satellite this tension has grown drastically. In this work we have explored the possible observational consequences of some of the most successful models that have been proposed to ease the Hubble tension, focusing specifically on new physics in the CMB damping tail.
@@
We find that some, but not all, of these models make predictions for clear deviations from $\Lambda$CDM in the small-scale CMB power spectrum. In particular, we find that WZDR (with or without interactions) predicts a decrement in power on small scales, with a possible upturn at even smaller scales. Pure EDE models all shows features in the TT and TE spectra around $\ell\sim 1000$ that might be borderline observable, as well as an overall upturn of the TT and EE power spectra at $\ell\sim 3000$. When a non-clustering dark matter component is added to EDE, as in the case of NEDE with a non-negligible trigger field energy density, the small scale power may be suppressed at $\ell\sim 3000$. Finally, we find that variations in the electron mass leads to a relatively flat fractional change in power.
@@
Near-future CMB data is expected to be able to potentially detect some, but not all, of the investigated solutions to the Hubble tension. The predicted deviations for WZDR are significantly larger than the expected error bars of near-future CMB experiments, indicating that its potential as a viable solution may be clearly decided relatively soon with the upcoming ACT DR6 and SPT datasets. On the other hand, EDE models (with or without a non-clustering matter component) predict deviations from $\Lambda$CDM at such small scales that one needs to be concerned about the accuracy of the modeling of the non-linear growth of structure as well as baryonic feedback. Once that modeling is under control it may be possible to detect these features with CMB observatories such as CMB-S4 and the Simons Observatory. Finally, variations in the electron mass make predictions for the small-scale CMB which may remain indistinguishable from $\Lambda$CDM even with future CMB observations in line with the analytical arguments of Schöneberg and Vacher (2025); Sekiguchi and Takahashi (2021).
@@
It is important to note that, in addition to having to develop new techniques to model the non-linear growth of structure, any future detection of these features also requires proper modeling of foreground contamination. For example, the upturn in the fractional power predicted by pure EDE models is at the several percent level – an effect that may be degenerate with terms in the foreground modeling.
@@
It is important to note that the predictions of these models appear in many other cosmological data sets. Shifts in the cosmological parameters for many of these models are likely to be observable in future late-time data such as from the large scale structure probed by galaxy surveys or from the luminosity distance of SNeIa.
@@
Note While this manuscript was being finalized the results of ACT DR6 were announced Louis et al. (2025); Calabrese et al. (2025); the main results were produced before the announcement. As anticipated by our analysis, ACT DR6 is unable to place significant constraints on EDE (see Sec. 5.2.1 and Fig. 12 of Ref. Calabrese et al. (2025)), and in fact shows a preference for EDE with an improvement in the $\chi^{2}$ of up to 6.6 (see Table 2). Similarly, Sec. 5.2.1 shows that when fit to ACT DR6, a varying electron mass plus curvature is able to raise the CMB-inferred value of $H_{0}$.
@@ Acknowledgements.
The authors thank Silvia Galli, Vivian Poulin, and Daniel Grin for useful conversations. TLS is supported by NSF Grants No. 2009377 and No. 2308173. Computational work for this was, in part, performed on Firebird, a cluster supported by Swarthmore College and Lafayette College. NS gratefully acknowledges the support by a Fraunhofer-Schwarzschild Fellowship at Universitäts-Sternwarte München.

### v2

@@ III.2 Early dark energy
We have computed the predictions from a scalar field EDE model following Refs. Poulin et al. (2019); Smith et al. (2020) using a scalar field potential which scales as $\phi^{6}$ around its minimum, thereby fixing the rate at which the EDE energy density redshifts after its peak to be $\rho_{\phi}\propto a^{-4.5}$ Turner (1983). In addition to this we have computed the predictions using a fluid parameterization Simon (2024); Lin et al. (2019); Poulin et al. (2023) in order to explore whether there are aspects of the small-scale CMB power spectrum in these models that are unique to the scalar field dynamics. In the fluid case we have left the equation of state after the peak EDE contribution as an additional free parameter.
The constraints for both these variations are shown in figure 2. The fluid approximation deviates slightly more than the axion-like early dark energy model. However, generally the two models behave very similarly. We do observe that they deviate especially in the temperature autocorrelation on the smallest scales (at $\ell\gtrsim 3000$), but in a way that is mostly compatible with the expected uncertainty of ACT DR6 and somewhat degenerate with the non-linear (baryonic feedback) treatment.
@@
We show the constraints for this model in figure 3. For the original cold NEDE model the constraints are almost identical to that of the axion-like EDE, except for a clear upturn in the TE spectra around $\ell\gtrsim 1500$ that could be detectable with near-future data. On the other hand, tNEDE clearly shows a very different phenomenology on the smallest scales, with a non-zero posterior for the energy density of the trigger field today, $\Omega_{0,t}=0.0042^{+0.0018}_{-0.0030}$ and a mass $m_{t}=2.16^{+0.45}_{-0.58}\times 10^{-27}\ {\rm eV}$. In this case the model is generally more consistent with $\Lambda$CDM due to the additional suppression of the small-scale power– $S_{8}=0.818^{+0.022}_{-0.019}$ in tNEDE versus $S_{8}=0.847\pm 0.013$ in NEDE. The two models show similar features in the TE spectrum around $\ell\sim 1000$ that might be detectable. We note that while the mean value in the temperature autocorrelation for this variation is much more discrepant from the $\Lambda$CDM reference due to the additional suppression of small-scale power, the $1\sigma$ contour is quite broad, making it somewhat uncertain how well this feature may be measured with future CMB data.
@@
In the absence of an ultra-light scalar field the predicted small-scale power spectra for all EDE models considered here are very similar, showing a clear upturn in both TT and EE starting around $\ell\gtrsim 3000$. As we argue in section IV, such a feature is generic for any model that addresses the Hubble tension by injecting a significant amount of energy density over a limited amount of time.
@@ III.3 Varying electron mass
There are several indications that our current standard model of particle physics is not the final description of the universe (lacking explanations for dark energy and dark matter, a neutrino mass mechanism, as well as a quantum theory of gravity) and it is natural to ask if fundamental ‘constants’ are allowed to vary in time. In particular, an electron mass, $m_{e}$, which was slightly larger around recombination than it is today has proven to be surprisingly effective at resolving the cosmic conundrum of the Hubble tension as well as being slightly favored by recent DESI data, see Hart and Chluba (2020); Sekiguchi and Takahashi (2021); Schöneberg et al. (2022); Schöneberg and Vacher (2025); Baryakhtar et al. (2024, 2025). With a larger value of $m_{e}$ the redshift of recombination is increased, leading to a smaller sound horizon and therefore a larger allowed Hubble constant while keeping the CMB angular sound horizon scale unchanged. Indeed, in Schöneberg and Vacher (2025); Sekiguchi and Takahashi (2021); Baryakhtar et al. (2024) it has been argued that the CMB might be entirely insensitive to such a change at first order if other cosmological parameters are varied concurrently. See Hart and Chluba (2018); Chiang and Slosar (2018); Hart and Chluba (2022); Chluba and Hart (2023); Hart and Chluba (2020); Sekiguchi and Takahashi (2021); Lynch et al. (2024); Baryakhtar et al. (2025) for further constraints, and Schöneberg and Vacher (2025) for a review.
In particular, the small-scale CMB is left nearly invariant if the relative abundance of baryons and photons, matter and radiation, and the angular size of the sound horizon at recombination is fixed Sekiguchi and Takahashi (2021); Baryakhtar et al. (2024). The resulting scaling gives, $h\propto m_{e}^{3.18}$ and $\Omega_{m}h^{2}\propto m_{e}$ Baryakhtar et al. (2024). The first scaling law tells us that to resolve the Hubble tension we need $m^{\rm early}_{e}/m^{\rm late}_{e}\simeq 1.02$ which, in turn, limits the physical matter density to increase by at most $\sim 2\%$. Large-scale CMB measurements and late-time probes of the expansion history (such as SNeIa and BAO) give tight constraints on $\Omega_{m}$, which, along with the required increase in $h$ (and, in turn, $\Omega_{m}h^{2}$), leads to limitations on this model’s ability to fully address the Hubble tension (even more complicated changes to the recombination history face the same constraints Lee et al. (2023)) [^2] As shown in the next section, dark radiation and EDE models are less subject to this since they introduce additional degeneracies allowing for $\Omega_{m}h^{2}$ to reach larger values.. One way to mitigate this is to include additional degrees of freedom for the late-time expansion history, such as allowing for non-zero curvature, as we do below.
Here we implement the model as described in Schöneberg and Vacher (2025) using a phenomenological approach – noting that a microphysical theory of electron mass variations is described in Ref. Baryakhtar et al. (2024) (see also Ref. Schöneberg and Vacher (2025)) and constrained, using a wide range of data sets, in Ref. Baryakhtar et al. (2025). In these theories the mass of the electron is controlled by an ultra-light scalar field which becomes dynamical some time between recombination and today.
@@
We describe the model using a single additional parameter $m_{e}^{\mathrm{early}}/m_{e}^{\mathrm{late}}$ describing the relative difference of the electron mass around CMB times compared to that measured in the laboratory today. We assume a transition in the dark ages where electromagnetic observational data is sparse, see also Ref. Schöneberg and Vacher (2025) for more details on other observational probes for this model, including a discussion on why BBN light element abundances are currently not constraining enough to restrict such variations.
@@
We show the predicted power spectra in figure 4. We observe that the varying $m_{e}$ model can leave the intermediate to small-scale CMB invariant with respect to $\Lambda$CDM and is therefore mostly indistinguishable from $\Lambda$CDM despite having a much larger value of $H_{0}$ (around $H_{0}=71.2\pm 0.7$km/s/Mpc with $m_{e}^{\mathrm{early}}/m_{e}^{\mathrm{late}}=1.0198\pm 0.0044$, a roughly 2% increase in the electron mass at CMB times) Sekiguchi and Takahashi (2021); Schöneberg and Vacher (2025); Baryakhtar et al. (2024). A a slightly large $H_{0}$ can be reached when including curvature (green curve and colored region in figure 4). In this case there are larger oscillations in the mean prediction, but overall the result is not far outside the expected uncertainty bands for near-future CMB surveys and the mean value is still mostly compatible with zero. In this case we find $H_{0}=71.8\pm 1.1$km/s/Mpc with $m_{e}^{\mathrm{early}}/m_{e}^{\mathrm{late}}=1.037\pm 0.013$, a roughly 4% increase in the electron mass at CMB times, and $\Omega_{k}=(-6.5\pm 4.5)\cdot 10^{-3}$.
@@ IV Discussion
One part of the explanation is due to the lack of balance between the small-scale effect of $\theta_{\rm d}$ and $n_{s}$. As was shown in Ref. (Smith and Poulin, 2024, III.A), both parameters need to increase to keep the small-scale CMB relatively unchanged (compared to the $\Lambda$CDM mean values). Their Eq. (7) reads
@@
Figure 7 shows the posteriors of $\theta_{\rm d}$ and $n_{s}$ for $\Lambda$CDM, WZDR, and axion-like EDE. The black line shows the degeneracy given in equation 1 with $\delta\theta_{\rm d}\equiv\theta_{\rm d}-\bar{\theta}^{\Lambda{\rm CDM}}_{\rm d}$ and $\delta n_{s}\equiv n_{s}-\bar{n}^{\Lambda{\rm CDM}}_{s}$. It is clear that the axion-like EDE model lies along this degeneracy whereas WZDR is significantly above it. Models which are above the line have a damping scale which is larger, leading to enhanced damping on small scales.
@@
Variations in the electron mass leave both the damping scale and $n_{s}$ unchanged as well as have tighter upper limits on $\Omega_{m}h^{2}$ due to the degeneracies introduced by these models between $m_{e}$, $H_{0}$ and $\Omega_{m}h^{2}$, as discussed in section III.3. Additionally, as argued in Schöneberg and Vacher (2025); Baryakhtar et al. (2024) there is a strong degeneracy for all underlying scales (damping, angular, equality, …) involved in the CMB that should cause this model to be very close to $\Lambda$CDM, which is what we observe.
@@ V Conclusions
With the advent of precision cosmology the Hubble tension has become a serious problem for the standard $\Lambda$CDM model. In particular, since the arcminute-scale CMB anisotropies have been observed by the Planck satellite this tension has grown drastically. In this work we have explored the possible observational consequences of some of the most successful models that have been proposed to ease the Hubble tension, focusing specifically on new physics in the CMB damping tail.
@@
We find that some, but not all, of these models make predictions for clear deviations from $\Lambda$CDM in the small-scale CMB power spectrum. In particular, we find that WZDR (with or without interactions) predicts a decrement in power on small scales, with an upturn at even smaller scales. Pure EDE models all show features in the TT and TE spectra around $\ell\sim 1000$ that might be borderline observable, as well as an overall upturn of the TT and EE power spectra at $\ell\sim 3000$. When a non-clustering dark matter component is added to EDE, as in the case of NEDE with a non-negligible trigger field energy density, the small scale power may be suppressed at $\ell\sim 3000$. Finally, we find that variations in the electron mass leads to a relatively flat fractional change in power.
Figure 8: The lensing potential power spectrum predictions from the best fit predictions of three of the models we consider here. Note that, as discussed in section IV the EDE prediction rises above the others at large multipoles.
Far future measurements of the lensing potential power spectrum may also play a role in distinguishing these models. We show a comparison between the best fit predictions of these models and our baseline $\Lambda$CDM model in figure 8. Limitations to measuring the lensing potential power spectrum at large angular scales likely prevents us from detecting a deficit of power there. However, the enhanced power in EDE models may be detected with far future CMB measurements, with the caveat that the way we model non-linear growth of structure will have a significant effect on these predictions.
It is also interesting to consider how ‘de-lensing’ may play a role in detecting these deviations Smith et al. (2012); Hotinli et al. (2022). However, we note that in the limit of reconstructing the true primary CMB some of the key features – such as the enhancement of power in pure EDE models – are removed (see figure 6). We leave an exploration of how de-lensing may affect our conclusions to future work.
@@
Near-future CMB data are expected to be able to potentially detect the damping tail predictions of some, but not all, of the investigated solutions to the Hubble tension. The predicted deviations for WZDR are significantly larger than the expected error bars of near-future CMB experiments, indicating that its potential as a viable solution may be clearly decided relatively soon with the upcoming ACT DR6 and SPT datasets. On the other hand, EDE models (with or without a non-clustering matter component) predict deviations from $\Lambda$CDM at such small scales that one needs to be concerned about the accuracy of the modeling of the non-linear growth of structure as well as baryonic feedback. Once that modeling is under control it may be possible to detect these features with CMB observatories such as CMB-S4 and the Simons Observatory. Finally, variations in the electron mass make predictions for the small-scale CMB which may remain indistinguishable from $\Lambda$CDM even with future CMB observations in line with the analytical arguments of Sekiguchi and Takahashi (2021); Schöneberg and Vacher (2025); Baryakhtar et al. (2024).
@@
In addition to having to develop new techniques to model the non-linear growth of structure, any future detection of these features also requires proper modeling of foreground contamination. For example, the upturn in the fractional power predicted by pure EDE models is at the several percent level – an effect that may be degenerate with terms in the foreground modeling.
This is not to say that improved CMB data cannot put pressure on these models. For example, improved measurements of the intermediate scale CMB ($\ell\sim 1000-2000$) have already put pressure on some of the models considered here. The most up-to-date analysis of the Planck satellite data (NPIPE) increased its precision relative to its original data release Aghanim et al. (2020b) (plik), leading to improved constraints on both pure EDE and NEDE models Efstathiou et al. (2024); Chatrchyan et al. (2025), with $H_{0}=71.81_{-0.87}^{+0.88}$ km/s/Mpc using plik becoming $H_{0}=71.25_{-0.81}^{+0.8}$ km/s/Mpc with NPIPE for the pure EDE model. On the other hand, constraints on the WZDR and varying electron mass models are only minimally changed with the updated Planck data: $H_{0}=71.76_{-0.82}^{+0.82}$ km/s/Mpc with plik becoming $H_{0}=71.83_{-0.83}^{+0.82}$ km/s/Mpc with NPIPE for WZDR and $H_{0}=72.18_{-0.9}^{+0.86}$ km/s/Mpc with plik going to $H_{0}=72.27_{-0.88}^{+0.88}$ km/s/Mpc with NPIPE for a varying electron mass with non-zero curvature. All of these constraints use the data sets described in section II.
@@
It is important to note that unique predictions of these models appear in many other cosmological data sets. Shifts in the cosmological parameters for many of these models are likely to be observable in future late-time data such as from the large scale structure probed by galaxy surveys or from the luminosity distance of SNeIa.
@@
Note While this manuscript was being finalized the results of ACT DR6 were announced Louis et al. (2025); Calabrese et al. (2025); the main results were produced before the announcement. As anticipated by our analysis, ACT DR6 is unable to place significant constraints on EDE (see Sec. 5.2.1 and Fig. 12 of Ref. Calabrese et al. (2025)), and in fact shows a slight preference for EDE with an improvement in the $\chi^{2}$ of up to 6.6 (see Table 2). Similarly, Sec. 5.2.1 shows that when fit to ACT DR6, a varying electron mass plus curvature is able to raise the CMB-inferred value of $H_{0}$.
@@ Acknowledgements.
The authors thank Silvia Galli, Vivian Poulin, Daniel Grin, Zachary J. Weiner, and Ali Rida Khalife for useful conversations. TLS is supported by NSF Grants No. 2009377 and No. 2308173. Computational work for this was, in part, performed on Firebird, a cluster supported by Swarthmore College and Lafayette College. NS gratefully acknowledges the support by a Fraunhofer-Schwarzschild Fellowship at Universitäts-Sternwarte München.